# CtRL-Sim: Reactive and Controllable Driving Agents with Offline Reinforcement Learning

### v2

@@ Abstract
Evaluating autonomous vehicle stacks (AVs) in simulation typically involves replaying driving logs from real-world recorded traffic. However, agents replayed from offline data are not reactive and hard to intuitively control. Existing approaches address these challenges by proposing methods that rely on heuristics or generative models of real-world data but these approaches either lack realism or necessitate costly iterative sampling procedures to control the generated behaviours. In this work, we take an alternative approach and propose CtRL-Sim, a method that leverages return-conditioned offline reinforcement learning to efficiently generate reactive and controllable traffic agents. Specifically, we process real-world driving data through a physics-enhanced Nocturne simulator to generate a diverse offline reinforcement learning dataset, annotated with various reward terms. With this dataset, we train a return-conditioned multi-agent behaviour model that allows for fine-grained manipulation of agent behaviours by modifying the desired returns for the various reward components. This capability enables the generation of a wide range of driving behaviours beyond the scope of the initial dataset, including adversarial behaviours. We demonstrate that CtRL-Sim can generate diverse and realistic safety-critical scenarios while providing fine-grained control over agent behaviours.
@@ 2.1 Our Approach to Controllable Simulation via Offline RL
We consider the common offline RL setup where we are given a dataset $\mathcal{D}$ of trajectories $\tau_{i}=\{\ldots,s_{t},a_{t},r_{t},\ldots\}$, with states $s_{t}\in\mathcal{S}$, actions $a_{t}\in\mathcal{A}$ and rewards $r_{t}$. These trajectories are generated using a (suboptimal) behaviour policy $\pi_{B}(a_{t}|s_{t})$ executed in a finite-horizon Markov decision process. The return at timestep $t$ is defined as the cumulative sum of scalar rewards obtained in the trajectory from timestep $t$, $G_{t}=\sum_{t^{\prime}=t}^{T}r_{t^{\prime}}$. The objective of offline RL is to learn policies that perform as well as or better than the best agent behaviours observed in $\mathcal{D}$.
@@ 3.1 Experimental Setup
Offline Reinforcement Learning Dataset To train our model, we curate an offline reinforcement learning dataset derived from the Waymo Open Motion dataset [14]. We first extend Nocturne by integrating a physics engine based on the Box2D library for enabling realistic vehicle dynamics and collisions, detailed in Appendix C.1. Each scene in the Waymo dataset is fed through the physics-enhanced Nocturne simulator to compute the per-timestep actions and factored rewards for each agent. Refer to Appendix C for more details regarding the offline RL dataset collection.
@@
We evaluate the following variants of the proposed CtRL-Sim model: 1. CtRL-Sim (Base) is the CtRL-Sim model trained on the offline RL dataset. 2. CtRL-Sim (No State Prediction) is the base model trained without the state prediction task. 3. CtRL-Sim (Positive Tilting) applies $\kappa_{c}=10$ tilting to all components $c$ of the base model. 4. CtRL-Sim (GT Initial Return) is similar to DT (GT Initial Return). For controllability evaluation, we evaluate on the CtRL-Sim base model and a finetuned CtRL-Sim model (CtRL-Sim FT). The finetuned model takes a trained base model and finetunes it on a dataset of simulated long-tail scenarios that we collect using an existing simulated collision generation method CAT [21]. This allows CtRL-Sim to be exposed to more long-tail collision scenarios during training, as the Waymo Open Motion dataset mainly contains nominal driving. We refer readers to Appendix H for details of CAT and our proposed finetuning procedure.
@@ 3.2 Results
In Table 1, we present the multi-agent imitation results comparing the CtRL-Sim model and its variants with imitation baselines. The CtRL-Sim models perform competitively with the imitation baselines, with the CtRL-Sim (Positive Tilting) model achieving a good balance between distributional realism (2nd in JSD), reconstruction performance (2nd in FDE, ADE), common sense (Tied 1st in Collision and Offroad Rate), and efficiency (5.4×5.4$\times$5.4 × faster than CTG++). Although DT (Max Return) attains equal collision and offroad rates as CtRL-Sim, this comes at the cost of substantially worse reconstruction performance. We further validate the importance of the future state prediction task, with CtRL-Sim (Base) outperforming CtRL-Sim (No State Prediction) across all metrics. The CtRL-Sim (Positive Tilting) model attains the best collision rate and offroad rate, demonstrating the effectiveness of exponential tilting for steering the model towards good driving behaviours.
@@
We emphasize that a distinctive feature of CtRL-Sim that distinguishes it from the imitation baselines in Table 1 is that it additionally enables intuitive control over the agent behaviours through exponential tilting of the return distribution. This contrasts with DT, which, although capable of generating suboptimal behaviours by specifying low initial return-to-gos, lacks intuitive control due to the prerequisite knowledge about the return-to-go values and an absence of an interpretable mechanism for behaviour modulation. By constrast, the exponential tilting of the predicted return distribution employed in CtRL-Sim has a clear interpretation: negative exponential tilting yields behaviours that are worse than the average behaviours learned from the dataset, while positive exponential tilting yields better-than-average behaviours. This provides a more intuitive interface to a practitioner who may aim to produce behaviours that are either less or more optimal than nominal driving behaviours.
@@
We show the results of our controllability evaluation in Figure 4. For each reward dimension $c$, we exponentially tilt $\kappa_{c}$ between -25 and 25 and observe how this affects the corresponding metric of interest. We also show the results of DT when conditioning on the minimum and maximum possible return. For both the base and finetuned CtRL-Sim models, we observe a relatively monotonic change in each metric of interest as the tilting coefficient is increased from -25 to 25. As the finetuned model is exposed to collision scenarios during finetuning, it demonstrates significant improvements over the base model in generating bad driving behaviours. Specifically, at -25 tilting, the finetuned model is able to generate 2.1×2.1$\times$2.1 × as many collisions and 1.8×1.8$\times$1.8 × as many offroad violations as the base model. Figure 5 (and 7, 8 in Appendix J) shows qualitatively the effects of tilting.
@@
Table 2 evaluates CtRL-Sim’s ability to produce adversarial agents that collide with a data-driven planner. We evaluate on a held-out test set of two-agent interactive scenarios from the Waymo dataset, where one interacting agent is controlled by the planner and the other is controlled by the adversary. We use a positively-tilted CtRL-Sim base model as our planner, due to its demonstrated ability to produce good driving behaviours in Table 1. For adversarial scenario generation, we compare the base CtRL-Sim model against the CtRL-Sim FT model. With -10 tilting applied, the finetuned model generates 238 more collisions with the planner than the base model over 1000 scenes, which we attribute to its exposure to simulated collision scenarios during finetuning. Notably, CtRL-Sim FT was finetuned in only 30 minutes on 1 NVIDIA A100-Large GPU and only 3500 CAT scenarios. This underscores CtRL-Sim’s capability to flexibly incorporate data from various sources through finetuning, thereby enabling the generation of new kinds of driving behaviours. Importantly, after finetuning, CtRL-Sim FT largely retains its ability to produce good driving behaviours. This is evidenced by a 21.1 percentage point decrease in the planner’s collision rate when using a +50 positively tilted finetuned model as the adversary.
In Table 2, we also compare CtRL-Sim against a state-of-the-art collision generation method, CAT [21], which uses a goal-conditioned trajectory forecasting model to select plausible adversarial trajectories that overlap with the ego plan. Although CAT generates more collisions with the planner, CAT is not controllable in that it cannot control the degree to which the agents are adversarial, which is a distinguishing feature of CtRL-Sim. Furthermore, CAT agents are non-reactive to the ego’s actions as the trajectory is fixed at the beginning of the simulation, which severely limits the realism of agents controlled by CAT. This is evidenced by a larger adversary collision speed than all finetuned CtRL-Sim models and is also validated qualitatively in the attached supplementary videos. As collision realism is hard to quantitatively assess, we further conduct a user study to confirm that CtRL-Sim adversarial scenarios are indeed more realistic than CAT adversarial scenarios, with details and results reported in Appendix I.
@@ 4 Related Work
Agent behaviour simulation involves modelling the behaviour of other agents in simulation, such as vehicles and pedestrians, to enable diverse and realistic interactions with the AV. Agent behaviour simulation methods can be categorized into rule-based and data-driven methods. Rule-based methods rely on human-specified rules to produce plausible agent behaviours, such as adhering strictly to the center of the lane [2, 3]. These methods often yield unrealistic and rigid agent behaviours that fail to capture the full spectrum of driving behaviours. Moreover, we are most interested in modelling long-tail behaviours, which are difficult to model with rules alone.
@@
To address these limitations, prior work has proposed learning generative models that aim to replicate agent behaviours found in real-world driving trajectory datasets [22, 23, 24, 4, 5, 25, 9]. These approaches draw significant inspiration from the extensive array of methods proposed for the task of joint motion prediction [26, 27, 28, 29, 30]; however, it’s crucial to distinguish that, unlike the open-loop nature of joint motion prediction, behaviour simulation operates in a closed-loop manner [31]. To improve the realism of the learned behaviours, other work has proposed using adversarial imitation learning [32] to minimize the behavioural discrepancy between expert and model rollouts [33, 34, 6] or RL to improve traffic rule compliance [35, 36]. While such methods demonstrate improved realism over rule-based methods, they lack the necessary control over the behaviours to enable the generation of targeted simulation scenarios for AV testing.
@@
More recent work has proposed more controllable behaviour simulation models by learning conditional models [10, 37, 7, 38, 8, 11, 12] that enable conditioning on a high-level latent variables [10, 37], route information [7], or differentiable constraints [8, 11, 39, 12, 40]. More recently, [41] used retrieval augmented generation to generate controllable traffic scenarios. However, these methods either lack interpretable control over the generated behaviours [37] or require costly test-time optimization procedures to steer the generated behaviours, such as latent variable optimization [10], Bayesian optimization [42, 43, 7], or the simulation of expensive diffusion processes [8, 44, 11, 39, 12, 40]. CtRL-Sim takes an alternative approach and learns a conditional multi-agent behaviour model that conditions on interpretable factorized returns. By exponentially tilting the predicted return distribution [16] at test time, CtRL-Sim enables efficient, interpretable, and fine-grained control over agent behaviours while being grounded in real-world data.
5 Conclusions
@@
We presented CtRL-Sim, a novel framework applying offline RL for controllable and reactive behaviour simulation. Our proposed multi-agent behaviour Transformer architecture allows CtRL-Sim to employ exponential tilting at test time to simulate a wide range of interesting agent behaviours. We present experiments showing the effectiveness of CtRL-Sim at producing controllable and reactive behaviours, while maintaining competitive performance on the imitation task compared to baselines. We hope CtRL-Sim can be further explored in future work to handle more reward function components, such as driving comfort and respecting traffic signalization, as well as explored in domains outside of autonomous driving.
@@ Appendix A Offline RL Approaches
Table 3 presents the different ways explored in the literature for learning policies in offline RL, and how these methods can sample return-maximizing actions at test time.
@@ C.2 Offline RL Dataset Collection
We compute rewards at each timestep, where our reward function is factored into three rewards components: a goal position reward, vehicle-vehicle collision reward, and vehicle-road-edge collision reward. The goal position reward is defined by:
@@ Appendix G Baseline Details
Actions Only
The actions-only baseline is encoder-decoder architecture implemented in exactly the same way as CtRL-Sim with a few ablations. These include removal of states and returns from the decoder sequence, and no state rollout predictions. This model was inspired by [9] but differs in that the model also has access to the agents’ goals.
Imitation Learning
The imitation learning baseline is also based on the CtRL-Sim multi-agent behaviour simulation architecture but lacks factorized return information. It is a step better than the actions-only baseline since it considers the states in the decoder sequence, and this is corroborated by its improved performance on multi-agent simulation results of Table 1.
DT
The Decision Transformer baseline is based on the seminal work [15]. We adopt an identical architecture to CtRL-Sim’s with some minor difference based on the algorithm. One such decision is the lack of a return prediction based on states, and instead returns are chosen at inference time based on domain knowledge. Returns are the first token fed to the decoder, followed by states in order to predict actions. We make the strong argument that this is suboptimal for controllability (results in Figure 4) and does not provide intuitive mechanism for selecting the return values to target.
CTG++
The CTG++ baseline is a diffusion model reimplementation of the recent work by [11]. We attempted to follow the architecture as closely as possible with a few minor differences. One such difference is that we diffuse over both states and actions, rather than diffusing over only actions and using an unicycle dynamics model to derive the states. We chose this approach because the underlying dynamics of the physics-enhanced Nocturne simulator is not necessarily governed by a unicycle dynamics model, and thus using a unicycle dynamics model would induce small errors in the derived states during training. We further note that we cannot replace the unicycle dynamics model with the Nocturne physics dynamics model as this forward model is not differentiable. We also condition on the present timestep and goal, to ensure fair comparison with CtRL-Sim. We note that at scene generation time, we diffuse over actions and states at a rate of 2 Hz which is consistent with the original CTG++ model. In addition, although we train with 100 diffusion steps, we run evaluations with 50 diffusion steps. As showing in Table 5, the difference in performance is insignificant. As we do not have information on the size of the network used in the original manuscript, we explored different hidden dimension sizes of the transformer architecture of the diffusion model, also shown in Table 5. A final difference is the future relative encoding. While CTG++ use the ground-truth to compute the relative encoding during training and a constant velocity model at test time, we opted to use the final historical timestep’s relative encoding. We found this approach to be more stable.

### v3

@@ Abstract
Evaluating autonomous vehicle stacks (AVs) in simulation typically involves replaying driving logs from real-world recorded traffic. However, agents replayed from offline data are not reactive and hard to intuitively control. Existing approaches address these challenges by proposing methods that rely on heuristics or generative models of real-world data but these approaches either lack realism or necessitate costly iterative sampling procedures to control the generated behaviours. In this work, we take an alternative approach and propose CtRL-Sim, a method that leverages return-conditioned offline reinforcement learning (RL) to efficiently generate reactive and controllable traffic agents. Specifically, we process real-world driving data through a physics-enhanced Nocturne simulator to generate a diverse offline RL dataset, annotated with various rewards. With this dataset, we train a return-conditioned multi-agent behaviour model that allows for fine-grained manipulation of agent behaviours by modifying the desired returns for the various reward components. This capability enables the generation of a wide range of driving behaviours beyond the scope of the initial dataset, including adversarial behaviours. We show that CtRL-Sim can generate realistic safety-critical scenarios while providing fine-grained control over agent behaviours.
@@ 2.1 Our Approach to Controllable Simulation via Offline RL
We consider the common offline RL setup where we are given a dataset $\mathcal{D}$ of trajectories $\tau_{i}=\{\ldots,s_{t},a_{t},r_{t},\ldots\}$, with states $s_{t}\in\mathcal{S}$, actions $a_{t}\in\mathcal{A}$ and rewards $r_{t}$. These trajectories are generated using a (suboptimal) behaviour policy $\pi_{B}(a_{t}|s_{t})$ executed in a finite-horizon Markov decision process. The return-to-go at timestep $t$ is defined as the cumulative sum of scalar rewards obtained in the trajectory from timestep $t$, $G_{t}=\sum_{t^{\prime}=t}^{T}r_{t^{\prime}}$. The objective of offline RL is to learn policies that perform as well as or better than the best agent behaviours observed in $\mathcal{D}$.
@@ 3.1 Experimental Setup
Offline RL Dataset We curate an offline RL dataset derived from the Waymo Open Motion Dataset (WOMD) [14]. We extend Nocturne by integrating a physics engine based on the Box2D library for enabling realistic vehicle dynamics and collisions, detailed in Appendix C.1. Each scene in the Waymo dataset is fed through the physics-enhanced Nocturne simulator to compute the per-timestep actions and factored rewards for each agent. The factored rewards comprise of a goal position, vehicle-vehicle collision, and a vehicle-road-edge collision reward, detailed in Appendix C.2.
@@
We evaluate the following variants of CtRL-Sim: 1. CtRL-Sim (Base) is the CtRL-Sim model trained on the offline RL dataset. 2. CtRL-Sim (No State Prediction) is the base model trained without the state prediction task. 3. CtRL-Sim (Positive Tilting) applies $\kappa_{c}=10$ tilting to all components $c$ of the base model. 4. Instead of predicting the return-to-go at each timestep, CtRL-Sim (GT Initial Return) uses the ground-truth initial return-to-go and, at each timestep, decrements the reward from the initial return-to-go until the episode terminates. For controllability evaluation, we evaluate on the base model and a finetuned CtRL-Sim model (CtRL-Sim FT). The finetuned model takes a trained base model and finetunes it on a dataset of simulated long-tail scenarios that we collect using an existing simulated collision generation method CAT [21]. This allows CtRL-Sim to be exposed to more long-tail collision scenarios during training, as the WOMD mainly contains nominal driving. We refer readers to Appendix H for details of CAT and our proposed finetuning procedure.
@@ 3.2 Results
Table 1 presents the multi-agent imitation results comparing the CtRL-Sim model and its variants with imitation baselines. The CtRL-Sim models perform competitively with the imitation baselines, with the CtRL-Sim (Positive Tilting) model achieving a good balance between distributional realism (2nd in JSD), reconstruction performance (2nd in FDE, ADE), common sense (Tied 1st in Collision and Offroad Rate), and efficiency (5.4×5.4$\times$5.4 × faster than CTG++). We note that while IL is faster than CtRL-Sim due to fewer tokens to decode and offers similar performance, IL is not controllable. Although DT (Max Return) attains equal collision and offroad rates as CtRL-Sim, the reconstruction performance is substantially worse. We further validate the importance of the future state prediction task, with CtRL-Sim (Base) outperforming CtRL-Sim (No State Prediction) across all metrics. The CtRL-Sim (Positive Tilting) model attains the best collision rate and offroad rate, demonstrating the effectiveness of exponential tilting for steering the model towards good driving behaviours.
@@
A distinctive feature of CtRL-Sim is that it enables intuitive control over the agent behaviours through exponential tilting of the return distribution. This contrasts with DT, which, although capable of generating suboptimal behaviours by specifying low initial return-to-gos, lacks intuitive control due to the prerequisite knowledge about the return-to-go values and an absence of an interpretable mechanism for behaviour modulation. By constrast, the exponential tilting employed in CtRL-Sim has a clear interpretation: negative exponential tilting yields behaviours that are worse than the average behaviours learned from the dataset, while positive exponential tilting yields better-than-average behaviours. We show the results of our controllability evaluation in Figure 4. For each reward dimension $c$, we exponentially tilt $\kappa_{c}$ between -25 and 25 and observe how this affects the corresponding metric of interest. We also show the results of DT when conditioning on the minimum and maximum possible return. For both the base and finetuned CtRL-Sim models, we observe a relatively monotonic change in each metric of interest as the tilting coefficient is increased. As the finetuned model is exposed to collision scenarios during finetuning, it demonstrates significant improvements over the base model in generating bad driving behaviours. Specifically, at -25 tilting, the finetuned model is able to generate 2.1×2.1$\times$2.1 × as many collisions and 1.8×1.8$\times$1.8 × as many offroad violations as the base model. Figure 5 (and 7, 8 in Appendix J) shows qualitatively the effects of tilting.
@@
Table 2 evaluates CtRL-Sim’s ability to produce adversarial agents that collide with a data-driven planner. We evaluate on a held-out test set of two-agent interactive scenarios from the Waymo interactive dataset, where one interacting agent is controlled by the adversary and the other is controlled by the planner. We use a positively-tilted CtRL-Sim base model as our planner, due to its demonstrated ability to produce good driving behaviours in Table 1. For adversarial scenario generation, we compare the base CtRL-Sim model against the CtRL-Sim FT model. With -10 tilting applied, the finetuned model generates 238 more collisions with the planner than the base model over 1000 scenes, which we attribute to its exposure to simulated collision scenarios during finetuning. Notably, CtRL-Sim FT was finetuned in only 30 minutes on 1 NVIDIA A100-Large GPU and only 3500 CAT scenarios. This underscores CtRL-Sim’s capability to flexibly incorporate data from various sources through finetuning, thereby enabling the generation of new kinds of driving behaviours. Importantly, after finetuning, CtRL-Sim FT largely retains its ability to produce good driving behaviours. This is evidenced by a 21.1 percentage point decrease in the planner’s collision rate when using a +50 positively tilted finetuned model as the adversary. We also compare CtRL-Sim against a state-of-the-art collision generation method, CAT [21], which uses a motion prediction model to select plausible adversarial trajectories that overlap with the ego plan. Although CAT generates more collisions, CAT is not controllable as it can’t control how adversarial the agents are, and CAT agents are non-reactive to the ego’s actions as the trajectory is fixed prior to the simulation, limiting its realism. This is evidenced by a larger adversary collision speed than all finetuned CtRL-Sim models and is also validated qualitatively in the supplementary video. We further conduct a user study to confirm that CtRL-Sim adversarial scenarios are indeed more realistic than CAT adversarial scenarios, with details and results reported in Appendix I.
@@ 4 Related Work
Agent behaviour simulation involves modelling the behaviour of other agents in simulation, such as vehicles and pedestrians. Agent behaviour simulation methods can be categorized into rule-based and data-driven methods. Rule-based methods rely on human-specified rules to produce plausible agent behaviours, such as adhering strictly to the center of the lane [2, 3]. These methods often yield unrealistic behaviours that fail to capture the full spectrum of driving behaviours. To address these limitations, prior work has proposed learning generative models that aim to replicate agent behaviours found in real-world driving trajectory datasets [22, 23, 24, 4, 5, 25, 9]. These approaches draw inspiration from methods for the task of joint motion prediction [26, 27, 28, 29, 30]; however, it’s crucial to distinguish that, unlike the open-loop nature of joint motion prediction, behaviour simulation operates closed-loop [31]. To improve the realism of the learned behaviours, other work has proposed using adversarial imitation learning [32] to minimize the behavioural discrepancy between expert and model rollouts [33, 34, 6] or RL to improve traffic rule compliance [35, 36]. While such methods demonstrate improved realism over rule-based methods, they lack the necessary control over the behaviours to enable the generation of targeted simulation scenarios for AV testing.
@@
More recent work has proposed more controllable behaviour simulation models by learning conditional models [10, 37, 7, 38, 8, 11, 12] that enable conditioning on a high-level latent variables [10, 37], route information [7], or differentiable constraints [8, 11, 39, 12, 40]. More recently, [41] used retrieval augmented generation to generate controllable traffic scenarios. However, these methods either lack interpretable control over the generated behaviours [37] or require costly test-time optimization procedures to steer the generated behaviours, such as latent variable optimization [10], Bayesian optimization [42, 43, 7], or the simulation of expensive diffusion processes [8, 44, 11, 39, 12, 40]. In contrast, CtRL-Sim offers a more efficient alternative and learns a conditional multi-agent behaviour model that conditions on interpretable factorized returns, thereby eliminating the need for costly test-time optimization. By exponentially tilting the predicted return distribution [16] at test time, CtRL-Sim enables efficient, interpretable, and fine-grained control over agent behaviours while being grounded in real-world data.
5 Conclusion
@@
We presented CtRL-Sim, a novel framework applying offline RL for controllable and reactive behaviour simulation. Our proposed multi-agent behaviour Transformer architecture allows CtRL-Sim to employ exponential tilting at test time to simulate a wide range of agent behaviours. We present experiments showing the effectiveness of CtRL-Sim at producing controllable and reactive behaviours, while maintaining competitive performance on the imitation task compared to baselines.
Limitations The learned policies produced by CtRL-Sim in its current form may make large-scale RL training on CtRL-Sim scenarios prohibitively expensive. However, we believe that further optimizations of the model and simulator such as exploring more compact models, utilizing model quantization or Flash Attention [45], reducing the planning frequency, or making the simulator GPU-accelerated can further improve the inference latency. We leave this for future work. Furthermore, CtRL-Sim enables control over the behaviours, but not the initial agent placements. We believe that a generative model that supports controlling both the initial agent placements and the agent behaviours is promising future work to explore.
@@ Appendix A Offline RL Approaches
We position CtRL-Sim within the field of offline RL. Table 3 presents the different ways explored in the literature for learning policies in offline RL, and how these methods can sample return-maximizing actions at test time. Prior methods in offline RL differ in how the policy is modelled during training and how inference is performed; we refer the reader to Table 3 in the Appendix A for a breakdown of the different approaches. As CtRL-Sim employs a return-conditioned policy for controllable simulation, we briefly present a class of related methods that learn return-conditioned policies (RCPs) [15, 16, 46, 47, 48, 18, 49] for offline RL. RCPs are concerned with learning the joint distribution of actions and returns, such that action sampling is conditioned on the return distribution. Instead of modelling the return distribution, the Decision Transformer [15] conditions the learned policy on the maximum observed return, $G_{\text{max}}$ in the dataset. That is, at inference time, the actions are sampled from $a_{t}\sim p_{\theta}(a_{t}|s_{t},G_{\text{max}})$. Lee et al. [16] instead propose to employ the learned return distribution, combined with exponential tilting, in order to sample high-return actions while remaining close to the empirical distribution. CtRL-Sim adopts exponential tilting of the predicted return distribution for each agent to finely control agent behaviour. Additionally, CtRL-Sim explicitly models the future sequences of states.
@@ C.2 Offline RL Dataset Collection
We compute rewards at each timestep, where our reward function is factored into three rewards components: a goal position reward, vehicle-vehicle collision reward, and vehicle-road-edge collision reward. We chose the reward functions based on the information provided by the Nocturne simulator, which includes indicators for goal success, vehicle-vehicle collisions, and vehicle-edge collisions. For each agent, their goal is set to the final state (i.e., position, heading, and velocity) of the ground-truth logged trajectory. The goal position reward is defined by:
@@ Appendix G Baseline Details
Actions Only The actions-only baseline is encoder-decoder architecture implemented in exactly the same way as CtRL-Sim with a few ablations. These include removal of states and returns from the decoder sequence, and no state rollout predictions. This model was inspired by [9] but differs in that the model also has access to the agents’ goals.
Imitation Learning The imitation learning baseline is also based on the CtRL-Sim multi-agent behaviour simulation architecture but lacks factorized return information. It is a step better than the actions-only baseline since it considers the states in the decoder sequence, and this is corroborated by its improved performance on multi-agent simulation results of Table 1.
DT The Decision Transformer baseline is based on the seminal work [15]. We adopt an identical architecture to CtRL-Sim’s with some minor difference based on the algorithm. One such decision is the lack of a return prediction based on states, and instead returns are chosen at inference time based on domain knowledge. Returns are the first token fed to the decoder, followed by states in order to predict actions. We make the strong argument that this is suboptimal for controllability (results in Figure 4) and does not provide intuitive mechanism for selecting the return values to target.
CTG++ The CTG++ baseline is a diffusion model reimplementation of the recent work by [11]. We attempted to follow the architecture as closely as possible with a few minor differences. One such difference is that we diffuse over both states and actions, rather than diffusing over only actions and using an unicycle dynamics model to derive the states. We chose this approach because the underlying dynamics of the physics-enhanced Nocturne simulator is not necessarily governed by a unicycle dynamics model, and thus using a unicycle dynamics model would induce small errors in the derived states during training. We further note that we cannot replace the unicycle dynamics model with the Nocturne physics dynamics model as this forward model is not differentiable. We also condition on the present timestep and goal, to ensure fair comparison with CtRL-Sim. We note that at scene generation time, we diffuse over actions and states at a rate of 2 Hz which is consistent with the original CTG++ model. In addition, although we train with 100 diffusion steps, we run evaluations with 50 diffusion steps. As showing in Table 5, the difference in performance is insignificant. As we do not have information on the size of the network used in the original manuscript, we explored different hidden dimension sizes of the transformer architecture of the diffusion model, also shown in Table 5. A final difference is the future relative encoding. While CTG++ use the ground-truth to compute the relative encoding during training and a constant velocity model at test time, we opted to use the final historical timestep’s relative encoding. We found this approach to be more stable.